OpenStudy (anonymous):

Hi, I'm confused why the answer to this question is y<7. Why wouldn't it be y<5? Here's the question: If y = 2x + 3 and x < 2, which of the following represents all the possible values for y?

OpenStudy (anonymous):

because... if y=2x+3 and x<2 now we have: let's take x=1 so 1<2 y=2*1+3 y=2+3 y=5 or x=0 so 0<2 now you have: y=2*0+3 y=3 each of them are true for y<7 because in the first one you have y=5 so 5<7 --True and the second one: y=3 so 3<7 ... hope this helps

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

My pleasure xD